# Pre-train, Align, and Disentangle: Empowering Sequential Recommendation with Large Language Models

### v1

@@ Abstract.
Sequential recommendation (SR) aims to model the sequential dependencies in users’ historical interactions to better capture their evolving interests. However, existing SR approaches primarily rely on collaborative data, which leads to limitations such as the cold-start problem and sub-optimal performance. Meanwhile, despite the success of large language models (LLMs), their application in industrial recommender systems is hindered by high inference latency, inability to capture all distribution statistics and catastrophic forgetting. To this end, we propose a novel Pre-train, Align, and Disentangle (PAD) paradigm to empower recommendation models with LLMs. Specifically, we first pre-train both the SR and LLM models to get collaborative and textual embeddings. Next, a characteristic recommendation-anchored alignment loss is proposed using multi-kernel maximum mean discrepancy with Gaussian kernels. Finally, a triple-experts architecture, consisting aligned and modality-specific experts with disentangled embeddings, is fine-tuned in a frequency-aware manner. Experiments conducted on three public datasets demonstrate the effectiveness of PAD, showing significant improvements and compatibility with various SR backbone models, especially on cold items. The implementation code and datasets will be publicly available.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
With the explosive growth of the interactions with web applications and platforms, research on sequential recommender(SR) system (Kang and McAuley, 2018a; Hidasi et al., 2015; Sun et al., 2019a) has garnered increasing attention. These models aim to capture the sequential dependencies in user behavior sequences, modeling both long-term and short-term preferences. However, most existing SR methods rely exclusively on tabular data or ID-based features as inputs, which often leads to challenges such as the cold-start problem (Sheng et al., 2024), resulting in suboptimal performance, particularly for less frequent users, items, and scenarios.
@@
To address the limitations of conventional sequential recommendation (SR) models, recent efforts have drawn inspiration from the success of large language models (LLMs) in understanding semantics and processing natural language. Several approaches have sought to enhance SR by leveraging LLM capabilities. On the one hand, methods like TALLRec (Bao et al., 2023) and LC-Rec (Zheng et al., 2024) explicitly adapt LLMs for recommendation by instruction tuning. They formulate the sequential recommendation task as text generation, i.e., predicting the title of the next item based on a user’s historical interactions. Afterward, LLaRA (Liao et al., 2023), and CoLLM (Zhang et al., 2023) also leverage LLM as recommender and they propose to concatenating the token embedding with collaborative embedding, which enables LLM to comprehend collaborative information.
@@
On the other hand, models such as CTRL (Li et al., 2023a) and Flip (Wang et al., 2023) propose aligning LLM embeddings with collaborative embeddings using contrastive learning, employing an InfoNCE alignment loss based on non-characteristic cosine or linear kernels. Besides, Taobao (Sheng et al., 2024) and AlignRec (Liu et al., 2024) propose to pre-train multi-modal representations and incorporates them into recommendation. Despite these advances, three practical challenges remain to be addressed:
@@
  • High Inference Latency of LLM-based Recommenders. LLM-based recommenders (Bao et al., 2023; Liao et al., 2023; Zhang et al., 2023) rely on the complex architecture of large language models to make predictions, which is impractical for real-world recommendation systems, where the inference of hundreds of items must be completed within hundreds of milliseconds (McMahan et al., 2013; He et al., 2014; Pan et al., 2024).
@@
  • Capturing All Statistics of Data Distribution. Many existing works employ non-characteristic kernels in their alignment loss (Li et al., 2023a; Wang et al., 2023). However, research on Reproducing Kernel Hilbert Space (Fukumizu et al., 2008; Muandet et al., 2017) has demonstrated that non-characteristic kernels fail to capture all statistical aspects of the data distribution, potentially leading to suboptimal performance. Evidence of the superiority of characteristic kernels in recommendation will be presented in Sec. 4.5.
@@
To address the aforementioned challenges, we propose a three-phase framework—Pre-Train, Align, and Disentangle—to empower sequential recommendation with LLMs. This framework consists of (1) pre-training LLM and recommendation models, (2) alignment between textual and collaborative with MK-MMD, and (3) supervised fine-tuning for recommendation. Specifically, during the pre-training phase, we employ a pre-trained LLM model, e.g., Llama3 (Dubey et al., 2024), alongside a sequential recommendation model, e.g., SASRec (Kang and McAuley, 2018a). The textual and collaborative embeddings are then obtained from these two models, respectively.
@@ 2.1. Problem Formulation
In our problem setting, we obtain a semantic domain $\mathcal{D}_{\text{text}}=\{(\{\mathbf{h}_{i}^{s}\},$ $\mathbf{x}_{i}^{s},y_{i})\}_{i=1}^{n}$ and a collaborative domain $\mathcal{D}_{\text{rec}}=\{(\{\mathbf{h}_{i}^{c}\},\mathbf{x}_{i}^{c},y_{i})\}_{i=1}^{n}$ with $n$ samples each. Specifically, $\{\mathbf{h}_{i}^{s}\}$, $\{\mathbf{h}_{i}^{c}\}$, $\mathbf{x}_{i}^{s}$, $\mathbf{x}_{i}^{c}$, and $y_{i}$ denotes behavioral item sequence in semantic embedding space, behavioral item sequence in collaborative embedding space, target item in semantic embedding space, target item in collaborative embedding space, and true label. The probability distributions characterized by these two domains are $P$ and $Q$.
@@ 4. Experiments
  • RQ1: How do PAD perform, compared with other SOTA LLM encoder methods?
@@ 4.1.3. Baselines
  • SASRec denotes the primary SASRec model, which is only trained by pure ID of items in the collaborative space.
@@
where $k_{u}\in\mathcal{K}$ in (3). We take $m=5$ and $\sigma=\{-3,-2,-1,0,1\}$ in the following experiments.
4.1.4. Implementation Details.
Our experiments are conducted on Tesla V100 GPUs and all the results shown are averaged over 3 runs. Detailed experimental settings are provided in Appendix. A.
@@ 4.2. Overall Performance (RQ1)
We further illustrate the effectiveness of our method on the cold items with the following analysis. Denote $\mathcal{P}_{\text{Top-10\% }}^{\text{ID}}$ and $\mathcal{P}_{\text{Bottom-10\% }}^{\text{ID}}$ as the group of item pairs with the top-10% and bottom-10% distance based on collaborative collaborative embeddings. We present the distribution of these item pairs under the distance distribution regarding the collaborative embeddings and textual embeddings in each model in Fig. 3. We observe that all existing methods, including SASRec, SMEM, and CTRL, can’t differentiate the distribution of $\mathcal{P}_{\text{Top-10\% }}^{\text{ID}}$ and $\mathcal{P}_{\text{Bottom-10\% }}^{\text{ID}}$ well. In contrast, our method succeeds in learning a generally larger textual distance for those item pairs with top collaborative distance and learning a smaller textual distance for those with bottom collaborative distance. We conclude with the following result:
@@ 4.5. Characteristic Alignment Loss (RQ3)
Their results are shown in Fig. 5(a), and we can see that characteristic kernels (i.e., Laplacian, and Gaussian) indeed perform better than non-characteristic ones. Notably, even if Info_NCE explicitly pulls positive samples closer and negative samples farther thus achieving better results than pure cosine kernel, it is still inferior to characteristic kernels because the mean embedding of its RKHS is not rich enough. We conclude:
@@ 4.7. Ablation Study (RQ5)
We conduct an ablation study on the MIND dataset to investigate the effect of each expert and the frequency-aware gating. The results are shown in Fig. 5(a). First, we remove each expert in the last phase and observe the performance deteriorates when we remove any of the three experts. Interestingly, removing the recommendation (ID) expert leads to the largest performance drop. This indicates that the alignment loss leads to inevitable information loss of the collaborative embeddings, even with our Rec-Anchored loss. Therefore, the recommendation-specific tower is essential to retain the collaborative signals.
@@
To compare the performance of LLM and pre-train language model (PLM) like BERT, we replace the LLM2Vec module with BERT-Base as encoder to obtain the text embedding of items. It achieves the performance of HR@10 = 18.3989 and nDCG@10 = 9.9810 on MIND, which is inferior to LLM2Vec achieving HR@10 = 18.6703 and nDCG@10 = 10.1515. We speculate the reasons are: 1) More parameters of Llama3-8B bring in more powerful encoding capability than BERT-Base with 110 million parameters only. 2) Compared with BERT, LLM2Vec is trained with more abundant text data on wiki corpus and adopts a different Unsupervised contrastive training (SimCSE) objective.
@@ 5.2. Large Language Model & Multi-Modal Rec
In the recommendation community, some works explicitly adopt alignment between tabular and textual data which could achieve information gain. For example, CTRL (Li et al., 2023a) leverages contrastive learning, and FLIP (Wang et al., 2023) uses data masking and reconstruction pre-training with contrastive learning. Afterward, some works like DisCo (Du et al., 2024) and DaRec (Yang et al., 2024) conduct disentanglement to capture the specific information. However, they simply separate the embeddings into different parts which may not be fully learned by the common single-expert structure. By contrast, our PAD adopts triple-experts in a frequency-aware manner to adaptively fuse different information, which achieves better disentanglement ability.
Meanwhile, numerous studies (Liu et al., 2019; Sun et al., 2019b; He and McAuley, 2016; Chen et al., 2017, 2019; Rajput et al., 2024; Singh et al., 2024; Zhang et al., 2024) have focused on enhancing recommender systems by incorporating multi-modal content. Recently, researchers have explored capturing user fine-grained preferences across different modalities. For instance, MMGCN (Wei et al., 2019) attempts to model user preferences using modal-specific user-item bipartite graphs.  (Bian et al., 2023) proposes a novel multi-modal mixture-of-experts (MoE) fusion network.  (Sheng et al., 2024) proposes a two-stage paradigm to enhance recommendation with multi-modality recommendation. M3CSR  (Chen et al., 2024) proposes a multi-modality framework for to tackle cold-start in short-video recommendation. Another line of research is to build semantic IDs based on LLM semantics (Rajput et al., 2024; Singh et al., 2024; Zhang et al., 2024). Existing methods either use a non-characteristic alignment loss or don’t employ a disentangled expert, while our method has both of these two critical parts.
@@ Appendix A EXPERIMENTAL SETTINGS
For LLM2Vec we choose the Llama3-8B as the base model with the unsupervised-trained LoRA weights. We leverage the news title on MIND and item title on Amazon to generate text embedding. Suppose $l$ is defined as the length of user’s historical interaction sequence. For all the three datasets, we select the users with no less than 5 interactions and extract the latest 23 items interacted. The task is to predict whether the user would click the $\left(l-2\right)$-th, $\left(l-1\right)$-th, and $l$-th item on training, validation, and test set, respectively, given the previous interacted items. Besides, AdamW (Loshchilov et al., [n. d.]) is adopted as optimizer. Batch size is 16, the early stop epoch is set to 10, and dropout probability is 0.1. $L$2 regularization is adopted with weight of 0.1.
@@ Appendix C Advanced Analysis
Denote $\mathcal{P}_{\text{Top-10\% }}^{\text{ID}}$ and $\mathcal{P}_{\text{Bottom-10\% }}^{\text{ID}}$ as the group of item pairs with the top-10% and bottom-10% distance based on collaborative ID embeddings. We present the distribution of these item pairs under the distance distribution regarding the collaborative and textual embeddings learned by each model on warm items in Fig. 7.
Specifically, we plot the collaborative and text embedding distance distribution of different methods including the original SASRec, PAD, SMEM, and CTRL. The green and red parts denote the bottom and top 10% distance in the collaborative space in the left figure, and we also plot the corresponding distance of these item pairs in the semantic space in the right figure. By comparing Fig. 3 and 7, it can be observed that the item pairs with close distance become more concordant in PAD, showing that PAD bring more alignment on cold items compared with other methods.

### v2

@@ Abstract.
Sequential Recommendation (SR) aims to leverage the sequential patterns in users’ historical interactions to accurately track their preferences. However, the primary reliance of existing SR methods on collaborative data results in challenges such as the cold-start problem and sub-optimal performance. Concurrently, despite the proven effectiveness of large language models (LLMs), their integration into commercial recommender systems is impeded by issues such as high inference latency, incomplete capture of all distribution statistics, and catastrophic forgetting. To address these issues, we introduce a novel Pre-train, Align, and Disentangle (PAD) framework to enhance SR models with LLMs. In particular, we initially pre-train both the SR and LLM models to obtain collaborative and textual embeddings. Subsequently, we propose a characteristic recommendation-anchored alignment loss using multi-kernel maximum mean discrepancy with Gaussian kernels. Lastly, a triple-experts architecture, comprising aligned and modality-specific experts with disentangled embeddings, is fine-tuned in a frequency-aware manner. Experimental results on three public datasets validate the efficacy of PAD, indicating substantial enhancements and compatibility with various SR backbone models, particularly for cold items. The code and datasets are accessible for reproduction [^1] https://github.com/Applied-Machine-Learning-Lab/PAD.
@@
🖂Corresponding author
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730059isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Recommender systems
@@ 1. Introduction
With the explosive growth of the interactions with web applications and platforms (Wang et al., 2023c, 2024b; Zhao et al., 2018a; Wang et al., 2024c), research on sequential recommender system (SRS) (Kang and McAuley, 2018a; Hidasi et al., 2015; Sun et al., 2019; Li et al., 2023c; Gao et al., 2024; Zhao et al., 2022) has garnered increasing attention. These models aim to capture the sequential dependencies in user behavior sequences (Zhang et al., 2024a, 2022), modeling both long-term and short-term preferences (Liu et al., 2024a). However, most existing SR methods rely exclusively on tabular data or ID-based features as inputs (Zhao et al., 2018b; Liu et al., 2023; Wang et al., 2023a), which often leads to challenges such as the cold-start problem (Sheng et al., 2024), resulting in suboptimal performance, particularly for less frequent users, items, and scenarios.
@@
To address the limitations of conventional sequential recommendation (SR) models, recent efforts have drawn inspiration from the success of large language models (LLMs) in understanding semantics and processing natural language (Xu et al., 2024). Several approaches have sought to enhance SR by leveraging LLM capabilities (Liu et al., 2024b). On the one hand, methods like TALLRec (Bao et al., 2023) and LC-Rec (Zheng et al., 2024) explicitly adapt LLMs for recommendation by instruction tuning. They formulate the sequential recommendation task as text generation, i.e., predicting the title of the next item based on a user’s historical interactions. Afterward, LLaRA (Liao et al., 2023), and CoLLM (Zhang et al., 2023) also leverage LLM as recommender and they propose to concatenating the token embedding with collaborative embedding, which enables LLM to comprehend collaborative information.
@@
On the other hand, models such as CTRL (Li et al., 2023a) and Flip (Wang et al., 2023b) propose aligning LLM embeddings with collaborative embeddings using contrastive learning, employing an InfoNCE alignment loss based on non-characteristic cosine or linear kernels. Besides, Taobao (Sheng et al., 2024) and AlignRec (Liu et al., 2024c) propose to pre-train multi-modal representations and incorporates them into recommendation. Despite these advances, three practical challenges remain to be addressed:
@@
  • High Inference Latency of LLM-based Recommenders. LLM-based recommenders (Bao et al., 2023; Liao et al., 2023; Zhang et al., 2023; Wang et al., 2024a) rely on the complex architecture of large language models to make predictions, which is impractical for real-world recommendation systems, where the inference of hundreds of items must be completed within hundreds of milliseconds (McMahan et al., 2013; He et al., 2014; Pan et al., 2024).
@@
  • Inability to Capture All Statistics of Data Distribution. Many existing works employ non-characteristic kernels in their alignment loss (Li et al., 2023a; Wang et al., 2023b). However, research on Reproducing Kernel Hilbert Space (Fukumizu et al., 2008; Muandet et al., 2017) has demonstrated that non-characteristic kernels fail to capture all statistical aspects of the data distribution, potentially leading to suboptimal performance. Evidence of the superiority of characteristic kernels in recommendation will be presented in Sec. 4.5.
@@
To address the aforementioned challenges, we propose a three-phase framework—Pre-Train, Align, and Disentangle—to empower sequential recommendation with LLMs. This framework consists of (1) pre-training LLM and recommendation models, (2) alignment between textual and collaborative with MK-MMD, and (3) supervised fine-tuning for recommendation. Specifically, during the pre-training phase, we employ a pre-trained LLM, e.g., Llama3 (Dubey et al., 2024), alongside a sequential recommendation model, e.g., SASRec (Kang and McAuley, 2018a). The textual and collaborative embeddings are then obtained from these two models, respectively.
@@ 2.1. Problem Formulation
First we provide the formulation and notations. In our problem setting, we obtain a semantic domain $\mathcal{D}_{\text{text}}=\{(\{\mathbf{h}_{i}^{s}\},$ $\mathbf{x}_{i}^{s},y_{i})\}_{i=1}^{n}$ and a collaborative domain $\mathcal{D}_{\text{rec}}=\{(\{\mathbf{h}_{i}^{c}\},\mathbf{x}_{i}^{c},y_{i})\}_{i=1}^{n}$ with $n$ samples each. Specifically, $\{\mathbf{h}_{i}^{s}\}$, $\{\mathbf{h}_{i}^{c}\}$, $\mathbf{x}_{i}^{s}$, $\mathbf{x}_{i}^{c}$, and $y_{i}$ denotes behavioral item sequence in semantic embedding space, behavioral item sequence in collaborative embedding space, target item in semantic embedding space, target item in collaborative embedding space, and true label. The probability distributions characterized by these two domains are $P$ and $Q$.
@@ 4. Experiments
  • RQ1: How does PAD perform, compared with other SOTA baseline methods?
@@ 4.1.3. Baselines
  • SASRec (Kang and McAuley, 2018b) denotes the primary SASRec model, which is only trained by pure ID of items in the collaborative space.
@@
where $k_{u}\in\mathcal{K}$ in (3). We take $m=5$ and $\sigma=\{-3,-2,-1,0,1\}$ in the following experiments. Besides, our experiments are conducted on Tesla V100 GPUs and all the results shown are averaged over 3 runs. Detailed experimental settings are provided in Appendix. A.
@@ 4.2. Overall Performance (RQ1)
We further illustrate the effectiveness of our method on the cold items with the following analysis. Denote $\mathcal{P}_{\text{Top-10\% }}^{\text{ID}}$ and $\mathcal{P}_{\text{Bottom-10\% }}^{\text{ID}}$ as the group of item pairs with the top-10% and bottom-10% distance based on collaborative collaborative embeddings. We present the distribution of these item pairs under the distance distribution regarding the collaborative embeddings and textual embeddings in each model in Fig. 3. We observe that all existing methods, including SASRec, SMEM, and CTRL, can not differentiate the distribution of $\mathcal{P}_{\text{Top-10\% }}^{\text{ID}}$ and $\mathcal{P}_{\text{Bottom-10\% }}^{\text{ID}}$ well. In contrast, our method succeeds in learning a generally larger textual distance for those item pairs with top collaborative distance and learning a smaller textual distance for those with bottom collaborative distance. Consequently, we conclude with the following result:
@@ 4.5. Characteristic Alignment Loss (RQ3)
Their results are shown in Fig. 5(a), and we can see that characteristic kernels (i.e., Laplacian, and Gaussian) indeed perform better than non-characteristic ones. Notably, even if Info_NCE explicitly pulls positive samples closer and negative samples farther thus achieving better results than pure cosine kernel, it is still inferior to characteristic kernels because the mean embedding of its RKHS is not rich enough. Therefore, we conclude that:
@@ 4.7. Ablation Study (RQ5)
We conduct an ablation study on the MIND dataset to investigate the effect of each expert and the frequency-aware gating. The results are shown in Fig. 5(b). First, we remove each expert in the last phase and observe the performance deteriorates when we remove any of the three experts. Interestingly, removing the recommendation (ID) expert leads to the largest performance drop. This indicates that the alignment loss leads to inevitable information loss of the collaborative embeddings, even with our Rec-Anchored loss. Therefore, the recommendation-specific tower is essential to retain the collaborative signals.
@@
To compare the performance of LLM and pre-train language model (PLM) like BERT, we replace LLM2Vec with BERT-Base as the text encoder of items. It achieves the performance of HR@10 = 18.3989 and nDCG@10 = 9.9810 on MIND, which is inferior to LLM2Vec achieving HR@10 = 18.6703 and nDCG@10 = 10.1515. We speculate the reasons are: 1) More parameters of Llama3-8B bring in more powerful encoding capability than BERT-Base with 110 million parameters only. 2) Compared with BERT, LLM2Vec is trained with more abundant text data on wiki corpus and adopts a different Unsupervised contrastive training (SimCSE) objective.
@@ 5. Related Work
This section summarizes the related work on sequential recommendation, large language model, and multi-modal recommendation.
@@ 5.2. Large Language Model & Multi-Modal Rec
In the recommendation community, some works explicitly adopt alignment between tabular and textual data which could achieve information gain. For example, CTRL (Li et al., 2023a) leverages contrastive learning, and FLIP (Wang et al., 2023b) uses data masking and reconstruction pre-training with contrastive learning. Afterward, some works like DisCo (Du et al., 2024) and DaRec (Yang et al., 2024) conduct disentanglement to capture the specific information. However, they simply separate the embeddings into different parts which may not be fully learned by the common single-expert structure. By contrast, our PAD adopts triple-experts in a frequency-aware manner to adaptively fuse different information, which achieves better disentanglement ability.
Meanwhile, numerous studies (Liu et al., 2019; Sun et al., 2019; He and McAuley, 2016; Chen et al., 2017, 2019; Rajput et al., 2024; Singh et al., 2024; Zhang et al., 2024b; Liang et al., 2023) focused on enhancing recommender systems by incorporating multi-modal content. Recently, researchers have explored capturing user fine-grained preferences across different modalities. For instance, MMGCN (Wei et al., 2019) attempts to model user preferences using modal-specific user-item bipartite graphs. M3SRec (Bian et al., 2023) proposes a novel multi-modal mixture-of-experts (MoE) fusion network. Taobao  (Sheng et al., 2024) proposes a two-stage paradigm to enhance recommendation with multi-modality recommendation. M3CSR  (Chen et al., 2024) proposes a multi-modality framework for to tackle cold-start in short-video recommendation. However, the existing methods either use a non-characteristic alignment loss or don’t employ a disentangled expert, while our method has both of these two critical parts.
@@ 6. Conclusion
ACKNOWLEDGEMENT
This research was partially supported by Tencent Rhino-Bird Focused Research Program, Research Impact Fund (No.R1015-23), and Collaborative Research Fund (No.C1043-24GF).
@@ Appendix A Experimental Settings
For LLM2Vec we choose the Llama3-8B as the base model with the unsupervised-trained LoRA weights. We leverage the news title on MIND and item title on Amazon to generate text embedding. Suppose $l$ is defined as the length of user’s historical interaction sequence. For all the three datasets, we select the users with no less than 5 interactions and extract the latest 23 items interacted. The task is to predict whether the user would click the $\left(l-2\right)$-th, $\left(l-1\right)$-th, and $l$-th item on training, validation, and test set, respectively, given the previous interacted items. Besides, AdamW (Loshchilov et al., 2017) is adopted as optimizer. Batch size is 16, the early stop epoch is set to 10, and dropout probability is 0.1. $L$2 regularization is adopted with weight of 0.1. The weight $\gamma$ is searched and set to 0.2.
@@ Appendix C Advanced Analysis
Denote $\mathcal{P}_{\text{Top-10\% }}^{\text{ID}}$ and $\mathcal{P}_{\text{Bottom-10\% }}^{\text{ID}}$ as the group of item pairs with the top-10% and bottom-10% distance based on collaborative ID embeddings. Apart from the results on cold items in Fig. 3, we present the distribution of these item pairs under the distance distribution regarding the collaborative and textual embeddings learned by each model on warm items in Fig. 7. It indicates that the distribution of $\mathcal{P}_{\text{Top-10\% }}^{\text{ID}}$ and $\mathcal{P}_{\text{Bottom-10\% }}^{\text{ID}}$ can not be differentiated in textual modality probably because all these recommendation models mainly rely on collaborative information for warm items.